Sulfur dioxide and oxygen combine to form sulfur trioxide as shown below. 

2SO2(g) + O2(g) ⇔ 2SO3(g)

At 1 000 K, the value for Kp is 3.4. At one point in time, measurements were taken and the following partial pressures were measured: 

PSO2 = 0.48 atm

PO2 = 0.18 atm

PSO3 = 0.72 atm

Calculate QP


1
Expert's answer
2021-04-02T05:16:57-0400

Using the equilibrium expression, the reaction quotient (Qp) will be:

Qp = P(O2)*P^2(SO2)/P^2(SO3) = 0.18atm*(0.48atm)^2/(0.72atm)^2 = 0.08